OpenStudy (anonymous):

Find 3 consecutive odd integers such that when the 1st is subtracted from 4 times the 3rd, the result is 7.

OpenStudy (anonymous):

help?

OpenStudy (raden):

i think there is somethink wrong here.. for the sum the third first odd number : 1+3+5=9

OpenStudy (anonymous):

thats exactly how the problem is given to me

OpenStudy (raden):

oh sorry, i dont read the problem with true.. let the number are x, x+2, x+4

OpenStudy (anonymous):

so how would i right the equation to solve?

OpenStudy (anonymous):

@RadEn?

OpenStudy (raden):

sorry, my connection internet low so i cant continu it last time let the number are x, x+2, x+4 4(x+4) - x = 7 4x+16-x = 7 3x = 7-16 3x = -9 x= -3 so, the number are : -3, -1, 1